OpenStudy (anonymous):

Tiara wants to divide a square piece of paper into two equalivalent triangles. If the square measures 20 cm on each side, what will the third side of each triangle measure? A. 8.9 cm C. 40 cm D. 28.3 cm

OpenStudy (stormfire1):

Since you know that 2 two sides of the right triangles created will be 20 cm, you can simply use the Pythagorean theorem:\[a^2+b^2=c^2\]Solved for the third side you get:\[\sqrt{a^2+b^2}=c\]\[\sqrt{20^2+20^2}=28.3cm\]